• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name J. Shields

    Log entry time 20:42:15 on March 8,2012

    Entry number 364774

    keyword=[Target] Left rate increased to 1.4 kHz. Appears to be an expected result of beam movement.

    Planned change in beam position resulted in slightly higher rate on the left arm. At first we tried to evaporate some helium, but it turns out this was an expected result of the movement. Temperature is still okay.